# Giovanni Servedio1,2     Alessandro De Bellis1     Dario Di Palma1 Vito Walter Anelli1     Tommaso Di Noia1

### v1

@@ Abstract
Factual hallucinations are a major challenge for Large Language Models (LLMs). They undermine reliability and user trust by generating inaccurate or fabricated content. Recent studies suggest that when generating false statements, the internal states of LLMs encode information about truthfulness. However, these studies often rely on synthetic datasets that lack realism, which limits generalization when evaluating the factual accuracy of text generated by the model itself. In this paper, we challenge the findings of previous work by investigating truthfulness encoding capabilities, leading to the generation of a more realistic and challenging dataset. Specifically, we extend previous work by introducing: (1) a strategy for sampling plausible true-false factoid sentences from tabular data and (2) a procedure for generating realistic, LLM-dependent true-false datasets from Question Answering collections. Our analysis of two open-source LLMs reveals that while the findings from previous studies are partially validated, generalization to LLM-generated datasets remains challenging. This study lays the groundwork for future research on factuality in LLMs and offers practical guidelines for more effective evaluation. Code is provided at our Github Repo.
@@
Giovanni Servedio1,2     Alessandro De Bellis1     Dario Di Palma1 Vito Walter Anelli1     Tommaso Di Noia1 1Politecnico di Bari, Italy     2Sapienza University of Rome, Italy name.surname@poliba.it
@@ 4 Novel Dataset Generation Strategies for Factuality Self-Evaluation
This section introduces two novel dataset generation strategies to investigate LLM factuality self-evaluation. To contextualize our approach, we first examine the limitations of the True-False dataset by Azaria and Mitchell (2023). While it provides a structured framework for evaluating LLMs, its construction imposes constraints that may limit generalizability of the findings. Specifically, true and false statements are derived from tabular data using predefined templates. We argue that this approch suffers from several limitations:
@@ 4.1 Perplexity-based Dataset Construction
Since perplexity can be interpreted as a measure of plausibility, we operate under the assumption that an LLM possesses factual knowledge about a fact if the fact is assigned a "sufficiently low perplexity". Conversely, a high perplexity score for the true statement suggests that the LLM lacks knowledge of the fact. Given the limitations discussed earlier, we aim to evaluate the LLM ability to discern between true and false statements when it possesses the relevant knowledge. Therefore, we exclude instances where the LLM exhibits limited knowledge about the true statement: if the true statement does not rank among the lowest $k$ perplexity candidates, the generation process for that instance is aborted, and the next true fact is considered. In practice, we define $k$ as $k=\alpha|C|$, where $\alpha$ is a hyperparameter ($10<\alpha<$ 1). This accounts for the varying cardinality of the property ranges in the dataset, ensuring that the threshold for "sufficiently low perplexity" is adjusted based on the number of possible values for a given property.
@@
The normalization guarantees that $P(c_{i})$ values are suitable for sampling. Finally, a mixture of top-k and nucleus sampling Holtzman et al. (2020) is employed to sample the candidate for insertion into the template. This process generates a coherent yet factually incorrect statement that is more realistic and closely aligned with the LLM internal token prediction patterns.
@@ 4.2 LLM-Generated Dataset Construction
The strategy in Section 4.1 constructs a balanced true-false dataset from tabular data but has inherent limitations. While enabling direct comparison with Azaria and Mitchell (2023), it restricts generative models to rigid templates, limiting expressiveness. Additionally, its reliance on fixed candidate sets can lead to easily classifiable false statements when the set is small. Finally, template-based sampling reduces diversity, likely due to token bias: we found LMs consistently assigning lower perplexity to certain property-object pairs, regardless of the subject. To more realistically assess an LLM factuality self-evaluation, we propose generating both true and false statements directly from the model, overcoming the limitations of template-based approaches. This involves (a) a method to elicit diverse factoid statements from the LLM, and (b) a strategy to annotate statement veracity, addressing biases and inconsistencies in the process. Consider a Question Answering dataset composed of $N$ questions, $D_{QA}=\{(q_{i},a_{i})\}_{i=1}^{N}$, where each question $q_{i}$ has a corresponding ground-truth answer $a_{i}$. Given a LLM $\mathcal{M}$, we prompt it with each question $K$ times, yielding a set of generated answers. This results in an extended dataset $D^{\mathcal{M}}_{QA}=\{(q_{i},a_{i},\{a^{\mathcal{M}}_{i,k}\}^{K}_{k=1})\}_{i=1}^{N}$. The LLM-generated answers in $D^{\mathcal{M}}_{QA}$ can be annotated using an oracle LLM, which we assume is able to evaluate the veracity of each answer $a^{\mathcal{M}}_{i,k}$ given the ground-truth answer $a_{i}$ and the question $q_{i}$. This operation results in an annotated dataset
@@
Increasing K enhances the reliability of responses by offering a more accurate evaluation of the LLM knowledge state regarding a question. This evaluation can lead to three possible outcomes: (i) a high proportion of correct answers suggests the LLM fully understands the facts; (ii) an high proportion of incorrect answers indicates the LLM lacks or has partial knowledge, preventing correct responses; (iii) a mix of correct and incorrect answers implies knowledge with a tendency toward hallucination. This study assumes that an LLM can only encode factuality regarding a generated fact if it has some knowledge about it. Therefore, we focus exclusively on the third scenario. This scenario also naturally provides balance, including true and false variations of the same fact. We define the correct answer ratio $p^{\mathcal{M}}_{i}=\frac{1}{K}\sum_{k=1}^{K}v^{\mathcal{M}}_{i,k}$. We consider questions whose $p_{i}$ is around 0.5 with a tolerance hyperparameter $\tau$, that is $|p^{\mathcal{M}}_{i}-0.5|<\tau$. The dataset is obtained by selecting the answers and their veracity labels satisfying the condition:
@@ 5.1 Dataset Generation Setup
To generate the datasets for training the probe and alleviate the limitations of Azaria and Mitchell (2023), we employ the perplexity-based sampling procedure described in Section 4.1. Specifically, we set the needed hyperparameters as follows: $\alpha=0.1$, $\beta=0.1$, $k=10$ and $p=0.9$ We do not sample a dataset for the Scientific Facts topic, as the original version was generated by ChatGPT and not from tabular data (additional details in Appendix B). Table 4 summarizes the statistics for the two refined True-False datasets, while Figure 1 shows the differences in average perplexity between the original and the proposed datasets.
@@
We construct an LLM-generated dataset following the procedure described in Section 4.2. Our experiments use TriviaQA Joshi et al. (2017), a dataset of question-answer pairs collected from 14 trivia and quiz-league websites. Given the limited computational resources, we limit our focus to the validation split of TriviaQA Wikipedia questions. For each LLM in our analysis, we set $K=10$ and $\tau=0.1$, and we refuse the answers composed of less than 5 tokens to improve self-consistency. For annotation, we use GPT-4o mini with a 3-shot learning strategy, where three examples are provided. More details are included in Appendix C.1, and Table 8 reports the dataset statistics.
@@ 5.3 Generalization of SAPLMA on LLM-generated Sentences
We extend our experiments with the SAPLMA classifier to a new setting, where both training and evaluation are conducted on sentences generated by an LLM and sourced from TriviaQA, leveraging the procedure described in Section 4.2. Table 6 reports the performance of SAPLMA on a set of factual statements extracted from TriviaQA. The results indicate that the currently used probes are inadequate for factuality self-assessment in real-world scenarios, as the observed accuracy does not reach a noteworthy threshold. Furthermore, following the suggestion of Azaria and Mitchell, we optimize the classification threshold; however, this adjustment yields no significant improvement in performance. For completeness, the interested reader may find the result obtained on the same dataset when training on the original True-False dataset of Azaria and Mitchell in Appendix D.
@@ 6 Related Work
Probing techniques have become central for the layer-wise interpretation of deep learning models Alain (2016). This approach was then extended to Large Language Models (LLMs) to assess LLMs ability to encode syntax and semantics Conneau et al. (2018); Tenney et al. (2019). Specifically, among the different applications, a technique that emerges is self-evaluation, defined as a model’s ability to assess the accuracy of its own outputs Kadavath et al. (2022). Among the various works, notable research by Kadavath et al. (2022) explored estimating a well-calibrated P(True) directly from output probabilities to reflect answer accuracy. Orgad et al. (2024) explore error detection and hallucination mitigation in Question Answering by probing the internal representations of question-answer pairs. Their study shows that while LLMs encode factuality signals, these signals do not generalize well across task-specific datasets, suggesting that factuality encoding is task-dependent rather than universal. While, Chen et al. (2024) introduce the INSIDE framework, which detects hallucinations using the EigenScore metric to measure self-consistency across multiple LLM outputs for a single input. Similarly, Zhang et al. (2024) leverage a probe model, PINOSE, trained via offline consistency-checking, to perform online hallucination detection. Although INSIDE and PINOSE primarily focus on self-consistency in internal representations across multiple generated responses, they do not interpret hidden states or look for a veracity encoding.
@@ 7 Conclusion and Future Work
There are a few aspects of our study that could be explored further. We made every effort to check the datasets for inconsistencies, but a more thorough manual verification by human annotators would be beneficial for ensuring their robustness and minimizing potential biases. Additionally, our analysis is based on a limited set of models. While these models provide valuable insights, it is possible that larger or more complex models could demonstrate enhanced performance, particularly in the context of self-evaluation. Future work could expand on this by incorporating a wider range of models to investigate whether scalability can lead to improved results.
@@
A major concern when working with LLMs is their tendency to generate factually inaccurate information. When training probe classifiers to assess factual accuracy, biases and beliefs from the LLM may transfer to the probe, potentially reinforcing cultural, demographic, or ideological biases in factuality self-evaluation. However, with careful design, probing techniques can be adapted not only to minimize bias but also to actively mitigate its consequences.
@@ A.2 LLM-Generated Trivia Facts Dataset
Table 8 illustrates the number of annotates facts extracted from TriviaQA Joshi et al. (2017) employing the procedure described in section 4.2. Statistics are presented for each of the 14 question sources.
@@ Appendix B Perplexity-Based Refinement of the True-False Dataset Azaria and Mitchell (2023)
We base our dataset generation strategy described in Section 4 on the same tabular data employed by Azaria and Mitchell, which was made available upon request. The properties used for sampling are analogous to the ones employed by the original authors in their dataset generation. However, for the Cities topic, we restrict the analysis to facts like "<city> is city in <country>." Differently from the original authors, we avoid generating facts such as "<city> is the name of a city/country" having only two possible values, potentially resulting in too many easy-to-classify samples.

### v3

# Giovanni Servedio1,2†     Alessandro De Bellis1†     Dario Di Palma1 Vito Walter Anelli1     Tommaso Di Noia1

@@ Abstract
Factual hallucinations are a major challenge for Large Language Models (LLMs). They undermine reliability and user trust by generating inaccurate or fabricated content. Recent studies suggest that when generating false statements, the internal states of LLMs encode information about truthfulness. However, these studies often rely on synthetic datasets that lack realism, which limits generalization when evaluating the factual accuracy of text generated by the model itself. In this paper, we challenge the findings of previous work by investigating truthfulness encoding capabilities, leading to the generation of a more realistic and challenging dataset. Specifically, we extend previous work by introducing: (1) a strategy for sampling plausible true-false factoid sentences from tabular data and (2) a procedure for generating realistic, LLM-dependent true-false datasets from Question Answering collections. Our analysis of two open-source LLMs reveals that while the findings from previous studies are partially validated, generalization to LLM-generated datasets remains challenging. This study provides a foundation for future research on factuality in LLMs and offers practical guidelines for more effective evaluation. Code is provided at our GitHub Repository.
@@
Giovanni Servedio1,2†     Alessandro De Bellis1†     Dario Di Palma1 Vito Walter Anelli1     Tommaso Di Noia1 1Politecnico di Bari, Italy     2Sapienza University of Rome, Italy name.surname@poliba.it
footnotetext: Corresponding authors.
@@ 4 Novel Dataset Generation Strategies for Factuality Self-Evaluation
This section introduces two novel dataset generation strategies to investigate LLM factuality self-evaluation. To contextualize our approach, we first examine the limitations of the True-False dataset by Azaria and Mitchell (2023). While it provides a structured framework for evaluating LLMs, its construction imposes constraints that may limit generalizability of the findings. Specifically, true and false statements are derived from tabular data using predefined templates. We argue that this approach suffers from several limitations:
@@ 4.1 Perplexity-based Dataset Construction
Since perplexity can be interpreted as a measure of plausibility, we operate under the assumption that an LLM possesses factual knowledge about a fact if the fact is assigned a "sufficiently low perplexity". Conversely, a high perplexity score for the true statement suggests that the LLM lacks knowledge of the fact. Given the limitations discussed earlier, we aim to evaluate the LLM ability to discern between true and false statements when it possesses the relevant knowledge. Therefore, we exclude instances where the LLM exhibits limited knowledge about the true statement: if the true statement does not rank among the lowest $k$ perplexity candidates, the generation of that instance is discontinued, and the next true fact is considered. In practice, we define $k$ as $k=\alpha|C|$, where $\alpha$ is a hyperparameter ($10<\alpha<$ 1). This accounts for the varying cardinality of the property ranges in the dataset, ensuring that the threshold for "sufficiently low perplexity" is adjusted based on the number of possible values for a given property.
@@
The normalization guarantees that $P(c_{i})$ values are suitable for sampling. Finally, a mixture of top-k and nucleus sampling Holtzman et al. (2020) is employed to sample the candidate for insertion into the template. Specifically, we apply top-k and nucleus sampling sequentially: we select the top-k highest-scoring candidates and then refine this set by choosing the smallest subset whose cumulative probability reaches a predefined threshold. This process generates a coherent yet factually incorrect statement that is more realistic and closely aligned with the LLM internal token prediction patterns. In Section 5.1 we detail the selected values for the hyperparameters $\alpha$, $\beta$, $k$, and $p$.
@@ 4.2 LLM-Generated Dataset Construction
The strategy in Section 4.1 constructs a balanced true-false dataset from tabular data but has inherent limitations. While enabling direct comparison with Azaria and Mitchell (2023), it restricts generative models to rigid templates, limiting expressiveness. Additionally, its reliance on fixed candidate sets can lead to easily classifiable false statements when the set is small. Finally, template-based sampling reduces diversity, likely due to token bias: we found LMs consistently assigning lower perplexity to certain property-object pairs, regardless of the subject. To more realistically assess an LLM factuality self-evaluation, we propose generating both true and false statements directly from the model, overcoming the limitations of template-based approaches. This involves (a) a method to elicit diverse factoid statements from the LLM, and (b) a strategy to annotate statement veracity, addressing biases and inconsistencies in the process. Consider a Question Answering dataset composed of $N$ questions, $D_{QA}=\{(q_{i},a_{i})\}_{i=1}^{N}$, where each question $q_{i}$ has a corresponding ground-truth answer $a_{i}$. Given a LLM $\mathcal{M}$, we prompt it with each question $K$ times, yielding a set of generated answers. This results in an extended dataset $D^{\mathcal{M}}_{QA}=\{(q_{i},a_{i},\{a^{\mathcal{M}}_{i,k}\}^{K}_{k=1})\}_{i=1}^{N}$. Following LLM-as-judge Gu et al. (2024); Calderon et al. (2025) practices, the LLM-generated answers in $D^{\mathcal{M}}_{QA}$ can be annotated using an oracle LLM, which we assume is able to evaluate the veracity of each answer $a^{\mathcal{M}}_{i,k}$ given the ground-truth answer $a_{i}$ and the question $q_{i}$. This operation results in an annotated dataset
@@
Increasing K enhances the reliability of responses by offering a more accurate evaluation of the LLM knowledge state regarding a question. This evaluation can lead to three possible outcomes: (i) a high proportion of correct answers suggests the LLM fully understands the facts; (ii) a high proportion of incorrect answers indicates the LLM lacks or has partial knowledge, preventing correct responses; (iii) a mix of correct and incorrect answers implies knowledge with a tendency toward hallucination. This study assumes that an LLM can only encode factuality regarding a generated fact if it has some knowledge about it. Therefore, we focus exclusively on the third scenario. This scenario also naturally produces a balanced dataset, including true and false variations of the same fact. We define the correct answer ratio $p^{\mathcal{M}}_{i}=\frac{1}{K}\sum_{k=1}^{K}v^{\mathcal{M}}_{i,k}$. We consider questions whose $p_{i}$ is around 0.5 with a tolerance hyperparameter $\tau$, that is $|p^{\mathcal{M}}_{i}-0.5|<\tau$. The dataset is obtained by selecting the answers and their veracity labels satisfying the condition:
@@ 5.1 Dataset Generation Setup
To generate the datasets for training the probe and alleviate the limitations of Azaria and Mitchell (2023), we employ the perplexity-based sampling procedure described in Section 4.1. Specifically, we set the needed hyperparameters as follows: $\alpha=0.1$, $\beta=0.1$, $k=10$ and $p=0.9$. We excluded the Scientific Facts topic from our perplexity-based sampling procedure, as its original version was generated by ChatGPT and not from tabular data (additional details in Appendix B). Table 4 summarizes the statistics for the two refined True-False datasets, while Figure 1 shows the differences in average perplexity between the original and the proposed datasets.
@@
We construct an LLM-generated dataset following the procedure described in Section 4.2. Our experiments use TriviaQA Joshi et al. (2017), a dataset of question-answer pairs collected from 14 trivia and quiz-league websites. Given the limited computational resources, we limit our focus to the validation split of TriviaQA Wikipedia questions. We construct a True-False dataset from two additional QA datasets, SQuAD 2.0 Rajpurkar et al. (2018) and TruthfulQA  Lin et al. (2022), employing the same approach, which we describe in Appendix E. For each LLM in our analysis, we set $K=10$ and $\tau=0.1$, and we discard the answers composed of less than 5 tokens to improve self-consistency. For annotation, we use GPT-4o mini with a 3-shot learning strategy, where three examples are provided. More details are included in Appendix C.1, and Table 8 reports the dataset statistics.
@@ 5.3 Generalization of SAPLMA on LLM-generated Sentences
We extend our experiments with the SAPLMA classifier to a new setting, where both training and evaluation are conducted on sentences generated by an LLM and sourced from TriviaQA, leveraging the procedure described in Section 4.2. Table 6 reports the performance of SAPLMA on a set of factual statements extracted from TriviaQA. The results indicate that the currently used probes are inadequate for factuality self-assessment in real-world scenarios, as the observed accuracy does not reach a noteworthy threshold. Furthermore, following the suggestion of Azaria and Mitchell, we optimize the classification threshold; however, this adjustment yields no significant improvement in performance. For completeness, the interested reader may find the result obtained on the same dataset when training on the original True-False dataset of Azaria and Mitchell in Appendix D. In addition, we report results on SQuAD 2.0 and TruthfulQA in Appendix E, which align closely with the results observed on TriviaQA, further supporting the consistency of our findings.
@@ 6 Related Work
Probing techniques have become central for the layer-wise interpretation of deep learning models Alain (2016). This approach was then extended to Large Language Models (LLMs) to assess LLMs ability to encode syntax and semantics Conneau et al. (2018); Tenney et al. (2019). Specifically, among the different applications, a technique that emerges is self-evaluation, defined as a model’s ability to assess the accuracy of its own outputs Kadavath et al. (2022). Among the various works, notable research by Kadavath et al. (2022) explored estimating a well-calibrated P(True) directly from output probabilities to reflect answer accuracy. Azaria and Mitchell (2023) demonstrate that LLMs are capable of detecting false claims in synthetic true-false datasets. Building upon this initial study, several works Marks and Tegmark (2023); Bürger et al. (2024); Levinstein and Herrmann (2024) investigate the generalization capabilities of probes in detecting hallucinations in LLMs. While these studies offer valuable insights into how factual knowledge is encoded in LLMs, they stop short of examining whether LLMs can assess the factuality of their own generations. Instead, they primarily evaluate models on artificially constructed or perturbed datasets, rather than on claims produced by the LLMs themselves. Orgad et al. (2024) explore error detection and hallucination mitigation in Question Answering by probing the internal representations of question-answer pairs. Their study shows that while LLMs encode factuality signals, these signals do not generalize well across task-specific datasets, suggesting that factuality encoding is task-dependent rather than universal. Gekhman et al. (2025) show that LLMs can internally encode knowledge, detectable by probing the hidden states, yet still fail to express it in their generated output. Chen et al. (2024) introduce the INSIDE framework, which detects hallucinations using the EigenScore metric to measure self-consistency across multiple LLM outputs for a single input. Similarly, Zhang et al. (2024) leverage a probe model, PINOSE, trained via offline consistency-checking, to perform online hallucination detection. Although INSIDE and PINOSE primarily focus on self-consistency in internal representations across multiple generated responses, they do not interpret hidden states or look for explicit factuality encoding.
@@ 7 Conclusion and Future Work
There are a few aspects of our study that could be explored further. We made every effort to check the datasets for inconsistencies, but a more thorough manual verification by human annotators would be beneficial for ensuring their robustness and minimizing potential biases. Additionally, our analysis is based on a limited set of models. While these models provide valuable insights, it is possible that larger or more complex models could demonstrate enhanced performance, particularly in the context of self-evaluation. Future work could expand on this, incorporating a wider range of models to investigate whether scalability can improve results.
Lastly, we highlight important considerations regarding the use of perplexity as a proxy for plausibility in relation to model knowledge. Recent findings suggest that LLMs often assign similar perplexity scores to both seen and unseen sentences Duan et al. (2024), which may limit the reliability of perplexity in distinguishing between plausible and implausible content. This challenge intersects with broader issues of memorization and data provenance. While our work focuses on extending and refining prior approaches to factuality self-evaluation, we recognize that a rigorous assessment of perplexity effectiveness in this context is a critical avenue for future investigation. We therefore encourage further research into more robust and theoretically grounded measures of plausibility for LLM-generated content.
@@
A major concern when working with LLMs is their tendency to generate factually inaccurate information. When training probe classifiers to assess factual accuracy, biases and beliefs from the LLM may transfer to the probe, potentially reinforcing cultural, demographic, or ideological biases in factuality self-evaluation. With careful design, probing techniques can be adapted not only to minimize bias but also to actively mitigate its consequences.
Appendix
@@ A.2 LLM-Generated Trivia Facts Dataset
Table 8 illustrates the number of annotated facts extracted from TriviaQA Joshi et al. (2017) employing the procedure described in section 4.2. Statistics are presented for each of the 14 question sources. Table 9 reports examples of generated facts for the Llama 2-7b and OPT-6.7b models.
@@
Model Label Sentence
Llama 2-7b 1 Arthur was married to Guinevere.
1 Arthur’s most famous wife was Guinevere.
1 Guinevere was married to Arthur.
1 Guinevere’s husband was King Arthur.
0 Lancelot was married to Guinevere.
0 Sir Lancelot was married to Queen Guinevere.
0 Sir Leonne was married to Queen Guinevere.
OPT-6.7b 0 Anakin Skywalker is Darth Vader’s son.
0 Darth Vader’s son is Darth Vader.
1 Darth Vader’s son is Luke Skywalker, a member of the Rebel Alliance.
1 Darth Vader’s son is Luke Skywalker.
1 Darth Vader’s son is Luke.
0 Darth Vader’s son is known as Darth Vader.
1 Luke’s father is Darth Vader.
Table 9: Examples of true and false facts generated by Llama 2-7b and OPT-6.7b based on the questions in TriviaQA Joshi et al. (2017), obtained following the procedure detailed in Section 4.2.
@@ Appendix B Perplexity-Based Refinement of the True-False Dataset Azaria and Mitchell (2023)
We base our dataset generation strategy described in Section 4 on the same tabular data employed by Azaria and Mitchell, which was made available upon request. The properties used for sampling are analogous to the ones employed by the original authors in their dataset generation. However, for the Cities topic, we restrict the analysis to facts like "<city> is city in <country>." Differently from the original authors, we avoid generating facts such as "<city> is the name of a city/country" having only two possible values, potentially resulting in too many easy-to-classify samples. Table 10 reports examples of false sentences generated following the perplexity-based sampling procedure described in Section 4.1 and the Llama 2-7b model.
True Sentence Original Negative Generated Negative (Llama 2-7b)
The crocodile has a habitat of freshwater. The crocodile has a habitat of various. The crocodile has a habitat of grassland/savanna.
The zebra has distinctive black and white stripes, which may help deter flies and provide camouflage. The zebra is a fast swimmer and can maintain high speeds for extended periods of time. The zebra is the fastest land animal, reaching speeds up to 60-70 mph.
Tantalum has the symbol Ta. Tantalum has the symbol Cs. Tantalum has the symbol Tm.
Table 10: Examples of false sentences generated with the Llama 2-7b model, using the perplexity-based generation strategy described in Section 4.1.
@@ Appendix D Additional Experiments on the LLM-Generated Dataset
Appendix E Experiments on Additional LLM-Generated Datasets
We extended our LLM-generated dataset construction approach (Section 4.2) to TruthfulQA Lin et al. (2022) and SQuAD 2.0 Rajpurkar et al. (2018) to further support our findings.
E.1 LLM-Generated Facts from TruthfulQA
Model Total Samples True Samples (%)
Llama 2-7b 842 49.76%
OPT-6.7b 331 49.85%
Table 12: Summary of the dataset obtained by extracting factoid sentences from TruthfulQA Lin et al. (2022), following the procedure described in Section 4.2.
OPT-6.7b Llama 2-7b
Random Split 1 Random Split 2 Random Split 1 Random Split 2
last 0.487 0.488 0.491 0.504
28 0.477 0.495 0.498 0.506
24 0.478 0.513 0.507 0.507
20 0.512 0.535 0.518 0.509
16 0.493 0.492 0.520 0.514
Table 13: Performance of SAPLMA on a fact dataset generated from TruthfulQA. A 50-50 holdout strategy is adopted. Results are shown for the Llama 2-7b and OPT-6.7b models.
OPT-6.7b Llama 2-7b
Random Split 1 Random Split 2 Random Split 1 Random Split 2
last 0.473 0.560 0.518 0.539
28 0.515 0.566 0.520 0.514
24 0.503 0.476 0.518 0.501
20 0.455 0.542 0.499 0.525
16 0.479 0.500 0.485 0.530
Table 14: Accuracy values obtained training SAPLMA on the original True-False dataset and testing on our facts dataset generated from TruthfulQA. A 50-50 holdout strategy is adopted. Results are shown for the Llama 2-7b and OPT-6.7b models.
Following the same procedure detailed in Section  4.2, we extract true and false facts from TruthfulQA. Table 12 reports the statistics of the LLM-generated datasets constructed with Llama 2-7b and OPT-6.7b. Table 13 presents the performance of the SAPLMA classifier when both trained and tested on the TruthfulQA dataset. Additionally, Table 14 shows the results obtained by SAPLMA on TruthfulQA after being trained on the original True-False dataset. Since the dataset lacks a topic-based split, we adopt a 50-50 holdout strategy for evaluation. The results on the TruthfulQA dataset support our findings: the probe classifier fails to generalize effectively to LLM-generated datasets.
E.2 LLM-Generated Facts from SQuAD 2.0
Topic Model Sentences (%) True
Economic_inequality Llama 2-7b 224 50.45%
OPT-6.7b 135 50.37%
Immune_system Llama 2-7b 167 49.70%
OPT-6.7b 72 50.00%
Rhine Llama 2-7b 120 50.83%
OPT-6.7b 59 47.46%
Warsaw Llama 2-7b 140 51.43%
OPT-6.7b 102 50.00%
Yuan Dynasty Llama 2-7b 109 49.54%
OPT-6.7b 53 52.83%
Table 15: Summary of the dataset obtained by extracting factoid sentences from SQuAD 2.0 Rajpurkar et al. (2018), following the procedure described in Section 4.2.
Dataset last 28 24 20 16
Economic_inequality Llama 0.525 0.510 0.515 0.492 0.502
OPT 0.517 0.516 0.499 0.496 0.464
Immune_system Llama 0.486 0.490 0.503 0.502 0.547
OPT 0.436 0.439 0.401 0.438 0.435
Rhine Llama 0.525 0.520 0.515 0.524 0.526
OPT 0.484 0.485 0.522 0.547 0.521
Warsaw Llama 0.558 0.542 0.545 0.546 0.555
OPT 0.521 0.531 0.545 0.544 0.509
Yuan_dynasty Llama 0.509 0.522 0.514 0.529 0.544
OPT 0.544 0.560 0.512 0.542 0.512
Table 16: Performance of SAPLMA on a fact dataset generated from SQuAD 2.0, restricted to the validation set and the top-5 most popular topics. The original topic-wise leave-one-out strategy is adopted. Results are shown for the Llama 2-7b and OPT-6.7b models.
last 28 24 20 16
Economic_inequality Llama 0.558 0.531 0.589 0.531 0.545
OPT 0.600 0.600 0.548 0.533 0.533
Immune_system Llama 0.569 0.593 0.593 0.575 0.599
OPT 0.611 0.611 0.611 0.708 0.542
Rhine Llama 0.467 0.492 0.475 0.442 0.525
OPT 0.525 0.559 0.593 0.593 0.576
Warsaw Llama 0.614 0.579 0.607 0.600 0.593
OPT 0.48 0.500 0.539 0.598 0.549
Yuan_dynasty Llama 0.523 0.523 0.541 0.505 0.550
OPT 0.453 0.547 0.491 0.528 0.547
Table 17: Accuracy values obtained training SAPLMA on the original True-False dataset and testing on the facts dataset generated from SQuAD 2.0. Results are shown for the Llama 2-7b and OPT-6.7b models.
For the SQuAD 2.0 dataset, we limit our dataset construction procedure (Section 4.2) to the validation split and the five most common topics, as the large number of questions makes applying our procedure to the full dataset computationally impractical due to time constraints. Table 16 presents the performance of the SAPLMA classifier when trained and tested on the SQuAD 2.0 dataset. Table 17, instead, reports the results obtained on SQuAD 2.0 after training on the original True-False dataset from Azaria and Mitchell (2023).
The results on the SQuAD 2.0 dataset further support our conclusion: existing probes fall short in accurately assessing factuality in real-world settings. Across all cases, accuracy remains below a meaningful threshold, highlighting the limited reliability of these methods for self-assessment in scenarios involving LLM-generated sentences.
Appendix F Examples of SAPLMA Predictions
The tables in this appendix provide illustrative examples that complement the quantitative results discussed in the main text. Specifically, Table 18 presents sentence-level predictions on facts generated from the TriviaQA dataset using Llama 2-7b. In this case, the probe classifier is trained adopting a topic-wise leave-one-out strategy, as in Table 6. Table 19 shows additional predictions obtained using the probe classifier, which is trained on all but one topic and tested on our refined version of the held-out topic (as in the Orig. setting of Table 5).
The code is provided in our GitHub Repository.
Sentence Predicted GT Confidence
Sir John Suckling was the first poet to be buried at Poet’s Corner in London’s Westminster Abbey. False False 0.373
The first poet to be buried at Poet’s Corner in London’s Westminster Abbey was Geoffrey Chaucer. True True 0.551
The poet who was the first to be buried in Westminster Abbey was Geoffrey Chaucer. False True 0.456
Roquefort cheese is made from sheep’s milk. True True 0.794
Roquefort cheese is made from cow’s milk. True False 0.851
The milk of sheep is used to make ’Roquefort Cheese’. False True 0.488
South Korea held its first ’Grand Prix’ motor race in 1999. False False 0.291
South Korea held its first ’Grand Prix’ motor race in 2010. False True 0.305
South Korea hosted its first ’Grand Prix’ motor race in 1966. False False 0.182
Table 18: Examples of sentence classification on facts generated from the TriviaQA dataset using Llama 2-7b. The original topic-wise leave-one-out strategy is adopted. Correct predictions are highlighted in green, while incorrect ones are shown in red.
Sentence Predicted GT Confidence
Tin has the symbol Sn. True True 0.990
Mercury is a liquid at room temperature and used in thermometers and some electrical switches. False True 0.113
Fluorine is in the Alkaline earth metal group. False False 0.012
Tellurium has the atomic number of 82. True False 0.998
Table 19: Examples of predictions using SAPLMA and Llama 2-7b. In these examples, the probe classifier is trained on the original version of the True-False dataset Azaria and Mitchell (2023) (excluding the element topic) and tested on our refined version of the element topic. Correct predictions are highlighted in green, while incorrect ones are shown in red.